### v2

# Approximating Submodular Matroid-Constrained Partitioning

Abstract: The submodular partitioning problem asks to minimize, over all partitions $\mathcal{P}$ of a ground set $V$, the sum of a given submodular function $f$ over the parts of $\mathcal{P}$. The problem has seen considerable work in approximability, as it encompasses multiterminal cuts on graphs, $k$-cuts on hypergraphs, and elementary linear algebra problems such as matrix multiway partitioning. This research has been divided between the fixed terminal setting, where we are given a set of terminals that must be separated by $\mathcal{P}$, and the global setting, where the only constraint is the size of the partition. We investigate a generalization that unifies these two settings: minimum submodular matroid-constrained partition. In this problem, we are additionally given a matroid over the ground set and seek to find a partition $\mathcal{P}$ in which there exists some basis that is separated by $\mathcal{P}$. We explore the approximability of this problem and its variants, reaching the state of the art for the special case of symmetric submodular functions, and provide results for monotone and general submodular functions as well.
Keywords: Approximation algorithms, Matroid constraints, Multiway cut, Submodular partitioning
1 Introduction
Many motivations, as well as techniques, for submodular optimization stem from graph partitioning problems, making them a natural starting point. Given an undirected graph $G=(V,E)$, edge weights $w\in\mathbb{R}^{E}_{+},$ and a set $T\subseteq V$ of $k$ terminals, the Multiway Cut (Multiway-Cut) problem asks for a partition $\mathcal{P}=(V_{1},\dots,V_{k})$ of $V$ minimizing $\sum_{i=1}^{k}d_{w}(V_{i})$ such that each part $V_{i}$ contains exactly one terminal from $T$. A well-known special case is $k=2$, which corresponds to the polynomial-time solvable Min-cut problem. For $k\geq 3$, the problem is NP-hard [8] but admits approximation algorithms [15]. In contrast, the $k$-cut problem ($k$-Cut) takes as input only the graph $G=(V,E)$, edge weights $w\in\mathbb{R}^{E}_{+}$, and an integer $k$, and asks for a partition $\mathcal{P}=(V_{1},\dots,V_{k})$ of $V$ into $k$ non-empty parts minimizing $\sum_{i=1}^{k}d_{w}(V_{i})$. This problem is polynomial-time solvable for every fixed $k$, but NP-hard when $k$ is part of the input. In this case, a $(2-2/k)$-approximation is possible [14], and this approximation is tight under the Small Set Expansion Hypothesis [11].
These problems have been generalized to arbitrary submodular functions. A set function $f\colon 2^{V}\to\mathbb{R}$ is submodular if $f(A)+f(B)\geq f(A\cup B)+f(A\cap B)$ for all $A,B\subseteq V$. Additionally, it is symmetric if $f(A)=f(V\setminus A)$ for all $A\subseteq V$, and monotone if $f(A)\leq f(B)$ for all $A\subseteq B\subseteq V$. For ease of discussion, we assume throughout that $f(\emptyset)=0$. In the Submodular $k$-partition Problem (Sub-k-P), we are given a submodular function $f\colon 2^{V}\to\mathbb{R}_{\geq 0}$ and an integer $k$, and the goal is to find a partition $\mathcal{P}=\{V_{1},\dots,V_{k}\}$ of $V$ into $k$ non-empty parts minimizing $\sum_{i=1}^{k}f(V_{i})$. When $f$ is the cut function of an undirected graph, then it is symmetric and submodular, hence we get back the $k$-Cut problem. Sub-k-P does not admit a constant-factor approximation assuming the Exponential Time Hypothesis (ETH) [7], and the best known approximation factor is $O(k)$ [17, 12]. However, for certain special classes of submodular functions, it admits constant factor approximations. For symmetric submodular functions (SymSub-k-P), the problem admits a tight $(2-2/k)$-approximation [17, 4, 13], while for monotone submodular functions (MonSub-k-P), the problem admits a tight $4/3$-approximation  [4, 13].
In the Submodular Multiway Partition problem (Sub-MP), we are given a submodular function $f\colon 2^{V}\to\mathbb{R}_{+}$ and $k$ terminals $t_{1},\dots,t_{k}$, and the goal is to find a partition $\mathcal{P}=(V_{1},\dots,V_{k})$ with $t_{i}\in V_{i}$ for every $i\in[k]$ minimizing $\sum_{i=1}^{k}f(V_{i})$. When $f$ is the cut function of an undirected graph, then we get back the Multiway-Cut problem. Sub-MP admits a $2$-approximation [5]. For symmetric submodular functions (SymSub-MP), the problem admits a $3/2$-approximation [5], while for monotone submodular functions (MonSub-MP), the problem admits a $4/3$-approximation. An important special case of MonSub-MP is when $f$ is a graph coverage function (GCov-MP), that is, $f(S)\coloneqq\sum_{uv\in E\colon\{u,v\}\cap S\neq\emptyset}w(uv)$ for every $S\subseteq V$. We note that the objective function of GCov-MP differs from the objective function of Multiway-Cut by an additive factor of $|E|$. Yet, GCov-MP differs substantially from Multiway-Cut in terms of approximability: GCov-MP admits a $1.125$-approximation [2], while the best known approximation for Multiway-Cut is $1.2965$ [15].
Algorithmic techniques for global and fixed terminal partitioning problems differ significantly: the key difference is that fixed terminal versions allow for linear programming techniques. On the other hand, approximation algorithms for global partitioning problems have relied on global structural aspects of submodular functions, such as Gomory-Hu trees, greedy splitting, and cheapest singleton algorithms. Nevertheless, the approximation results are not always so different. For example, both MonSub-k-P and MonSub-MP admit a $4/3$-approximation. Previous work has also generalized these two regimes in some way, such as in [1] and [6]. One of the motivations of this work is to present a general model for partitioning problems that unifies the global and fixed-terminal variants, and to investigate approximation algorithms for this general model. The choice of encoding the terminal choices as the bases of some matroid is not without precedent: the same was effective for the minimum Steiner tree problem [3].
1.1 Results
We introduce a new problem that unifies the global and the fixed terminal partitioning problems. The Submodular Matroid-constrained Partitioning problem (Sub-MCP) is defined as follows:
Sub-MCP
Input: A submodular function $f\colon 2^{V}\to\mathbb{R}_{+}$ given by a value oracle, and a rank-$k$ matroid $\mathcal{M}=(V,\mathcal{B})$ given by an independence oracle, where $\mathcal{B}$ is the family of bases of $\mathcal{M}$.
Goal: Find a solution to $\min\left\{\sum_{i=1}^{k}f(V_{i})\colon\{V_{i}\}_{i=1}^{k}\text{ is a partition of }V\text{ s.t. there is }B\in\mathcal{B}\text{ with }|B\cap V_{i}|=1\text{ for all }i\in[k]\right\}.$
In other words, our aim is to find a $k$-partition that separates the elements of some basis of the matroid. If the input function is symmetric or monotone submodular, then we denote the resulting problem as SymSub-MCP or MonSub-MCP, respectively. A special case of SymSub-MCP is Matroid-constrained Multiway Cut (MC-Multiway-Cut), formally defined as follows.
MC-Multiway-Cut
Input: A graph $G=(V,E)$, edge weights $w\in\mathbb{R}^{E}_{+}$, and a rank-$k$ matroid $\mathcal{M}=(V,\mathcal{B})$ given by an independence oracle, where $\mathcal{B}$ is the family of bases of $\mathcal{M}$.
Goal: Find a solution to $\min\left\{\sum_{i=1}^{k}d_{w}(V_{i})\colon\{V_{i}\}_{i=1}^{k}\text{ is a partition of }V\text{ s.t. there is }B\in\mathcal{B}\text{ with }|B\cap V_{i}|=1\text{ for }i\in[k]\right\}.$
We study several algorithmic approaches for various special classes of matroids. Sub-MCP contains Hypergraph $k$-cut as a special case, so it does not admit a constant-factor approximation assuming the ETH. The first problem we study is SymSub-MCP. In this case, many of the algorithms for Sub-k-P extend to Sub-MCP while achieving the same approximation guarantees.
Algorithm 1 constructs a Gomory-Hu tree for a given symmetric submodular function and then greedily selects the minimum-weight edges for which a basis exists that respects the partition induced by the chosen edges.
Theorem 1
The Gomory-Hu tree algorithm provides a $(2-2/k)$-approximation for SymSub-MCP.
The greedy splitting algorithm from [17] proceeds in a greedy fashion somewhat differently than the Gomory-Hu tree approach. It first finds the cheapest submodular cut $S$ of $V$, and then iteratively finds, for a given intermediate partition $\mathcal{P}_{i}$, the next cheapest $\operatorname*{arg\,min}\{f(X)+f(W\setminus X)-f(W):X\subset W,W\in\mathcal{P}_{i}\}$ split to increase the size of the partition by one. This algorithm, as we show in Lemma 15, generalizes to the matroid-constrained versions as well to get the same $(2-2/k)$-approximation factor.
Theorem 2
The greedy splitting algorithm provides a $(2-2/k)$-approximation for SymSub-MCP.
Next, we turn to monotone submodular $k$-partitioning. Here, a natural adaptation of the cheapest singleton algorithm is a fast $(2-1/k)$-approximation. In the problem without the matroid constraint, the cheapest singleton algorithm proceeds by choosing $k-1$ elements with the cheapest singleton cost, namely such that $f(v_{1})\leq\dots\leq f(v_{k-1})\leq f(v_{j})$ for all $j\geq k$. It then returns the partition ($(\{v_{1}\},\dots,\{v_{k-1}\},\{v_{k},\dots,v_{n}\})$. To see why this yields a $(2-1/k)$-approximation, let $\mathcal{P}^{*}=\{V_{1}^{*},\dots,V_{k}^{*}\}$ be the optimal partition, ordered such that $f(V_{1}^{*})\leq\dots\leq f(V_{k}^{*})$. By monotonicity, we have $\sum_{i=1}^{k-1}f(v_{i})\leq\sum_{i=1}^{k-1}f(V_{i}^{*})\leq(1-1/k)\sum_{i=1}^{k}f(V_{i}^{*})$. Moreover, monotonicity and submodularity imply that $f(\{v_{k},\dots,v_{n}\})\leq f(V)\leq\sum_{i=1}^{k}f(V_{i}^{*})$. For the matroid-constrained setting, we can weight the matroid by the value of $f$ on the singletons. We then select a minimum weight independent set of size $k-1$ in the matroid, and return the partition where these $k-1$ elements are singleton classes, and the last class consists of the remaining elements. The approximation factor follows by the same argument.
We can obtain a slightly better bound using the greedy splitting technique from [17] by extending their main lemma to the matroid-constrained setting.
Theorem 3
The greedy splitting algorithm provides a $(2-2/k)$-approximation for MonSub-MCP.
We additionally show that this analysis is tight for the greedy splitting algorithm, even without the matroid constraints, i.e., there exist instances of MonSub-k-P for which the approximation factor of greedy splitting is at least $(2-2/k)$ – see Lemma 16.
Another consequence of the applicability of the greedy splitting algorithm is that we can extend the linear approximation guarantee from [17] to general Sub-MCP. We state this in Theorem 4.
Theorem 4
The greedy splitting algorithm provides a $(k-1)$-approximation for Sub-MCP.
Finally, we consider the special case of graph coverage functions. We denote by GCov-MCP the special case of Sub-MCP where the input function $f$ is a graph coverage function, defined as $f(S)\coloneqq\sum_{uv\in E\colon\{u,v\}\cap S\neq\emptyset}w(uv)$ for every $S\subseteq V$. Note that this function is monotone, and only differs from graph cut functions in terms of approximation, as $f(\mathcal{P})=d_{w}(\mathcal{P})+|E|$ for any partition $\mathcal{P}$, where $d_{w}(\mathcal{P})$ denotes the total weight of edges whose two endpoints are in diferent classes of $\mathcal{P}$. In this case, we can improve upon the result for monotone functions to achieve a factor that is identical to the current best factor for the global case. The proof is short and simple, so we include it here.
Theorem 5
The Gomory-Hu tree algorithm provides a $4/3$-approximation for GCov-MCP.
Proof: Let $\mathcal{P}$ be the partition returned by the Gomory-Hu tree algorithm and $\mathcal{P}^{*}$ be an optimal partition minimizing the graph coverage function $f$, which is also the partition minimizing $d_{w}$. Then, Theorem 1 shows $d_{w}(\mathcal{P})\leq(2-2/k)d_{w}(\mathcal{P}^{*})\leq 2d_{w}(\mathcal{P}^{*})$. Let $c=\frac{d_{w}(\mathcal{P}^{*})}{|E|}$. Then
$\frac{f(\mathcal{P})}{f(\mathcal{P}^{*})}\leq\frac{2d_{w}(\mathcal{P}^{*})+|E|}{d_{w}(\mathcal{P}^{*})+|E|}=\frac{2d_{w}(\mathcal{P}^{*})+d_{w}(\mathcal{P}^{*})/c}{d_{w}(\mathcal{P}^{*})+d_{w}(\mathcal{P}^{*})/c}=\frac{2c+1}{c+1}.$
On the other hand, $d_{w}(\mathcal{P}^{*})+|E|=(c+1)|E|$, and trivially $f(\mathcal{P})\leq 2|E|,$ so
$\frac{f(\mathcal{P})}{f(\mathcal{P}^{*})}\leq\frac{2|E|}{(c+1)|E|}=\frac{2}{c+1}.$
Thus $\frac{f(\mathcal{P})}{f(\mathcal{P}^{*})}\leq\max_{0\leq c\leq 1}\min\left\{\frac{2c+1}{c+1},\frac{2}{c+1}\right\}=4/3$, with the maximum attained at $c=1/2$.  $\Box$
Our results are summarized in Table 1.
Matroid-Constrained Global Fixed Terminals
General $O(k)$-apx (Thm. 4) $O(k)$-apx [12, 17] $2$-apx [5]
Symmetric $(2-2/k)$-apx (Thms. 1, 2) $(2-2/k)$-apx [17] $(3/2)$-apx [5]
Monotone $(2-2/k)$-apx (Thm. 3) $(4/3)$-apx [4] $(4/3)$-apx [2]
Graph Coverage $(4/3)$-apx (Thm. 5) $(4/3)$-apx [4] $(9/8)$-apx [2]
Table 1: Comparison of our approximation results (first column) with the current best for global and fixed-terminal settings across different classes of submodular functions.
2 Preliminaries
Notation.
We denote by $\mathbb{R}_{+}$ and $\mathbb{Z}_{+}$ the sets of nonnegative reals and integers, respectively. For a positive integer $k$, we use $[k]\coloneqq\{1,\dots,k\}$. For a set function $f:2^{V}\to\mathbb{R}_{+}$ and a partition $\mathcal{P}$ of $V$, we use the notation $f(\mathcal{P})\coloneqq\sum_{X\in\mathcal{P}}f(X)$. For a graph $G=(V,E)$ and $X\subseteq V$, let $\delta_{G}(X)$ denote the set of edges leaving $X$, that is, $\delta_{G}(X)\coloneqq\{uv\in E\colon|\{u,v\}\cap X|=1\}$. The cut function of $G$ is then defined as $d_{G}(X)\coloneqq|\delta_{G}(X)|$ for $X\subseteq V$. We omit the subscript $G$ when the graph is clear from the context. If edge weights $w\in\mathbb{R}^{E}_{+}$ are given, we use $d_{w}(X)\coloneqq\sum_{e\in\delta(X)}w(e)$. In directed graphs, we use $\delta^{in},d^{in}$, and $\delta^{out},d^{out}$ for incoming and outgoing edges, respectively. For a graph $G=(V,E)$ and a partition $\mathcal{P}=\{V_{1},\dots,V_{k}\}$ of $V$, $\delta_{G}(\mathcal{P})\coloneqq\cup_{i=1}^{k}\delta_{G}(V_{i})$ is called the boundary of $\mathcal{P}$. If edge weights $w\in\mathbb{R}^{E}_{+}$ are also given, we use $d_{w}(\mathcal{P})=w(\delta_{G}(\mathcal{P}))$.
For a subset $F\subseteq E$, we denote the graph obtained by deleting the edges in $F$ by $G-F$. When inserting (deleting) a single element $e$ from a set $X$, we often use $X+e$ $(X-e)$ for $X\cup\{e\}$ $(X\setminus\{e\})$.
Matroids.
A matroid $\mathcal{M}=(V,\mathcal{I})$ is defined by its ground set $V$ and its family of independent sets $\mathcal{I}\subseteq 2^{V}$ that satisfies the independence axioms: (I1) $\emptyset\in\mathcal{I}$, (I2) $X\subseteq Y,\ Y\in\mathcal{I}\Rightarrow X\in\mathcal{I}$, and (I3) $X,Y\in\mathcal{I},\ |X|<|Y|\Rightarrow\exists e\in Y-X\ s.t.\ X+e\in\mathcal{I}$. The subsets of $S$ not in $\mathcal{I}$ are called dependent. The rank $r(X)$ of a set $X$ is the maximum size of an independent set in $X$. The rank of the matroid is $r(V)$. The maximal independent subsets of $V$ are called bases and their family is usually denoted by $\mathcal{B}$. The matroid is uniquely defined by its rank function or by its family of bases, so we also use the notation $\mathcal{M}=(V,r)$ or $\mathcal{M}=(V,\mathcal{B})$. An inclusionwise minimal non-independent set is a circuit.
Given a matroid $\mathcal{M}=(V,\mathcal{I})$, its $k$-truncation is the matroid $(V,\mathcal{I}_{k})$ with $\mathcal{I}_{k}=\{X\in\mathcal{I}\mid|X|\leq k\}$. We denote the $k$-truncation of $\mathcal{M}$ by $(\mathcal{M})_{k}$. For a set $Z\subseteq V$, the contraction of $Z$ results in a matroid $\mathcal{M}/Z=(V\setminus Z,\mathcal{I}^{\prime})$ in which a set $X\subseteq V\setminus Z$ is independent if and only if $X\cup I$ is independent for any maximum independent set of $\mathcal{M}$ in $Z$. Let $D=(V,A)$ be a directed graph, $V_{1},V_{2}\subseteq V$, and $M_{1}=(V_{1},\mathcal{I}_{1})$ be a matroid. We say that a set $Y\subseteq V_{2}$ is linked to another set $X\subseteq V_{1}$ if $|X|=|Y|$ and there exists $|X|$ pairwise vertex-disjoint directed paths from $X$ to $Y$. Then the matroid $M_{2}=(V_{2},\mathcal{I}_{2})$ defined by $\mathcal{I}_{2}\coloneqq\{I_{2}\subseteq V_{2}\colon\text{$I_{2}$ is linked to$I_{1}$ for some $I_{1}\in\mathcal{I}_{1}$}\}$ is called the matroid of $M_{1}$ induced by $D$ on $V_{2}$.
A laminar matroid is a matroid $\mathcal{M}=(V,\mathcal{I})$ where =$\mathcal{I}=\{F\subseteq V\colon|F\cap V_{i}|\leq g_{i}\ \text{for each $i\in[q]$}\}$ for some laminar family $V_{1},\dots,V_{q}$ of subsets of $V$ and upper bounds $g_{1},\dots,g_{q}\in\mathbb{Z}_{+}$. A rank-$r$ matroid $\mathcal{M}=(V,\mathcal{I})$ is paving if every set of size at most $r-1$ is independent, or equivalently, if every circuit has size at least $r$. Such matroids have a nice characterization using hypergraphs [10, 16].
Proposition 6
For a non-negative integer $r$, a ground set $V$ of size at least $r$, and a (possibly empty) family =$\mathcal{H}=\{H_{1},\dots,H_{q}\}$ of proper subsets of $V$ such that $|H_{i}\cap H_{j}|\leq r-2$ for $1\leq i<j\leq q$, the family =$\mathcal{B}_{\mathcal{H}}=\{X\subseteq V\mid|X|=r,\ X\not\subseteq H_{i}\ \text{for }i=1,\dots,q\}$ forms the set of bases of a paving matroid. Furthermore, every paving matroid can be obtained in this way.
As is common in matroid algorithms, we assume that matroids are given by an independence oracle, and we measure the running time in terms of the number of oracle calls and other elementary operations. For simplicity, we refer to an algorithm as “polynomial time” if its number of oracle calls and elementary operations is polynomial in the size of the ground set.
3 Symmetric Submodular Functions
We prove Theorem 1 in this section. The proof builds on a method outlined in [1]. First, we show that MC-Multiway-Cut is solvable on trees by a type of greedy algorithm. Then we use this on the Gomory-Hu tree of a symmetric submodular function to obtain an approximation algorithm for SymSub-MCP.
3.1 MC-Multiway-Cut on Trees
As a first step, we show that MC-Multiway-Cut is solvable in polynomial time on trees. Interestingly, in this case the problem reduces to finding a minimum weight basis of a matroid. Recall that in this setting, we are given a tree $G=(V,E)$ together with edge weights $w\in\mathbb{R}^{E}_{+}$ and a matroid $\mathcal{M}=(V,\mathcal{I})$. Let us define the following family of subsets of $E$:
(1) =$\mathcal{I}^{\prime}=\{X^{\prime}\subseteq E\colon\text{the components $V_{1},\dots,V_{|X^{\prime}|+1}$ of $G-X^{\prime}$ have a transversal independent in $\mathcal{M}$}\}.$
In other words, an edge set $X^{\prime}$ is in $\mathcal{I}^{\prime}$ if there exists $v_{i}\in V_{i}$ for every connected component $V_{i}$ of $G-X^{\prime}$ such that $\{v_{1},\dots,v_{|X^{\prime}|+1}\}$ forms an independent set of $\mathcal{M}$.
Lemma 7
If $\mathcal{M}$ has rank at least $1$, then $\mathcal{I}^{\prime}$ satisfies the independence axioms.
Proof: We verify the axioms one by one. Axiom (I1) clearly holds, since the rank of $\mathcal{M}$ is at least one, and any non-loop $v\in V$ is a transversal.
To see (I2), let $Y^{\prime}\in\mathcal{I}^{\prime}$ and $X^{\prime}\subseteq Y^{\prime}$. Let $\mathcal{V}$ and $\mathcal{U}$ be the families of connected components of $G-Y^{\prime}$ and $G-X^{\prime}$, respectively. Since $X^{\prime}\subseteq Y^{\prime}$, the partition $\mathcal{U}$ forms a coarsening of $\mathcal{V}$. By the definition of $\mathcal{I}^{\prime}$, $Y^{\prime}\in\mathcal{I}^{\prime}$ implies that there exists a transversal $Y$ of $\mathcal{V}$ that is independent in $\mathcal{M}$. Since $\mathcal{U}$ is a coarsening of $\mathcal{V}$, a properly chosen subset $X$ of $Y$ will form a transversal of $\mathcal{U}$, thus showing that $X^{\prime}\in\mathcal{I}^{\prime}$.
Finally, we verify (I3). Take any $X^{\prime},Y^{\prime}\in\mathcal{I}^{\prime}$ with $|X^{\prime}|<|Y^{\prime}|$. Let $X,Y\in\mathcal{I}$ be the corresponding independent sets in $\mathcal{M}$ chosen such that $|X\cap Y|$ is as large as possible. By axiom (I3) for $\mathcal{I}$, there exists $u\in Y\setminus X$ such that $X+u\in\mathcal{I}$. Let $C$ be the component of $G-X^{\prime}$ containing $u$, and $v$ be the element of $X$ in $C$. If $v\notin Y$, then $X-v+u\in\mathcal{I}$ is still a transversal of the conencted somponents of $G-X^{\prime}$, contradicting the choice of $X$ and $Y$. Thus $v\in Y$. As $u,v\in Y$, there exists an edge $e\in Y^{\prime}$ that lies on the path between $u$ and $v$ in $G$. Then $X^{\prime}+e$ is in $\mathcal{I}^{\prime}$ as desired, as it has $X+u$ as an independent set of representatives.  $\Box$
Given the independence oracle access to $\mathcal{M}$, we can construct an independence oracle to the matroid $\mathcal{M}^{\prime}=(E,\mathcal{I}^{\prime})$, as deciding whether there is an independent set of representatives for a given partition of $V$ reduces to the intersection of a partition matroid and $\mathcal{M}$.
Remark 8
It is worth mentioning that $\mathcal{I}^{\prime}$ can be identified with the independent set of a matroid of $\mathcal{M}$ induced by a directed graph. To see this, pick an arbitrary vertex $r\in V$, and orient the edges of $G$ such that the resulting directed graph is an in-arborescence rooted at $r$. Then, subdivide every arc $uv$ with a new vertex $x_{uv}$, resulting in arcs $ux_{uv}$ and $x_{uv}v$. Finally, add a new vertex $x$ to the graph together with the arc $rx$. Note that the vertex set of the directed graph thus obtained is $V\cup\{x_{uv}\colon uv\in E\}\cup\{x\}$. Now consider the matroid obtained by taking the matroid of $\mathcal{M}$ induced by the directed graph on $\{x_{uv}\colon uv\in E\}\cup\{x\}$, and contracting $x$ in it. It is not difficult to check that $Y\subseteq\{x_{uv}\colon uv\in E\}$ is independent in this matroid if and only if the corresponding subset of edges is in $\mathcal{I}^{\prime}$.
Corollary 9
MC-Multiway-Cut is solvable in polynomial time on trees.
Proof: First, we observe that MC-Multiway-Cut has an optimal solution $\mathcal{P}=\{V_{1},\dots,V_{k}\}$ where $G[V_{i}]$ is connected for every $i\in[k]$. To show this, we choose $\mathcal{P}$ so that $|\delta_{G}(\mathcal{P})|$ is smallest, and choose a basis $B$ with $|B\cap V_{i}|=1$ for all $i\in[k]$. Suppose for contradiction that some $G[V_{i}]$ is not connected, and let $U$ be a component that does not contain $B\cap V_{i}$. Since $G$ is connected, there exists an edge from $U$ to some $V_{j}$. The partition $\mathcal{P}^{\prime}=\mathcal{P}\setminus\{V_{i},V_{j}\}\cup\{V_{j}\cup U,V_{i}\setminus U\}$ is also optimal and $|\delta_{G}(\mathcal{P}^{\prime})|<|\delta_{G}(\mathcal{P})|$, a contradiction.
By the above observation and Lemma 7, the minimum weight matroid-constrained multiway cut on trees corresponds to the minimum weight basis of the matroid $\mathcal{M}^{\prime}=(E,\mathcal{I}^{\prime})$. Thus, MC-Multiway-Cut can be solved in polynomial time by using the greedy algorithm to find a minimum weight basis of $\mathcal{M}^{\prime}$.  $\Box$
3.2 Symmetric Submodular Functions
Goemans and Ramakrishnan [9] were the first to note that symmetric submodular systems have a Gomory-Hu tree. Given a symmetric submodular function $f$ over a ground set $V$, its Gomory-Hu tree is a tree $H=(V,F)$ with a weight function $w_{H}\in\mathbb{R}^{F}_{+}$ that encodes the minimum $s-t$ cuts for each pair $s,t$ of vertices in the following sense: given a path $P_{s,t}$ between $s$ and $t$, $\min_{e\in P_{s,t}}w_{H}(e)=\min_{s\in S\subseteq V-t}f(S)$. Furthermore, the two components of the tree obtained by removing the edge of minimum $w_{H}$-weight on the path give the two sides of a minimum $s-t$ cut in the submodular system $f$.
The greedy algorithm solves MC-Multiway-Cut in the special case when $G$ is a tree. The classical $(2-2/k)$-approximation for Multiway-Cut uses 2-way cuts coming from the Gomory-Hu tree, and so does the $(2-2/k)$-approximation for $k$-Cut. We follow a similar approach to obtain a $(2-2/k)$-approximation for SymSub-MCP, presented in Algorithm 1. The algorithm can be interpreted as taking the minimum edges in the Gomory-Hu tree as long as they allow a valid system of representatives.
Algorithm 1 Approximation algorithm for SymSub-MCP
1:Input: A symmetric, submodular function $f\colon 2^{V}\to\mathbb{R}_{+}$ and a rank-$k$ matroid $\mathcal{M}$.
2:Output: A feasible partition $\mathcal{P}$.
3:Compute the Gomory-Hu tree $H=(V,E)$ of $f$ and its weight function $w_{H}$.
4:Let $\mathcal{M}^{\prime}=(E,\mathcal{I}^{\prime})$ be the matroid defined as in (1) using $\mathcal{M}$ and $H$.
5:Set $C\leftarrow\emptyset$.
6:while $|C|<k-1$ do
7:     $e\leftarrow\operatorname*{arg\,min}\{w_{H}(e)\colon e\notin C,C+e\in\mathcal{I}^{\prime}\}$
8:     $C\leftarrow C+e$
9:end while
10:Return the connected components of $(V,E\setminus C)$.
Theorem 10
Algorithm 1 provides a $(2-2/k)$-approximation to SymSub-MCP.
[…]
Let us direct each $g_{i}$ away from $x$ in the tree $H$, and let $\vec{g_{i}}$ denote the directed edge. We then have
[…]
$\leq\sum_{i=1}^{k}\left(\sum_{\vec{g_{j}}\in\delta^{out}(V_{i})}f(U(g_{j}))+\sum_{\vec{g_{j}}\in\delta^{in}(V_{i})}f(V\setminus U(g_{j}))\right)=\sum_{i=1}^{k}\sum_{g_{j}\in\delta(V_{i})}w_{H}(g_{j})=\sum_{i=1}^{k}d_{w_{H}}(V_{i}),$
where the inequality follows by submodularity. Then
[…]
where the second inequality holds because $\{V_{1},\dots,V_{k}\}$ is an optimal solution of MC-Multiway-Cut on $H$ by Lemma 7, and the second to last inequality holds because of the assumption that $V^{*}_{k}$ has maximum $f$ value among the classes. This concludes the proof of the theorem.  $\Box$
3.3 Intersection of Two Matroids
Lemma 7 implies the solvability on trees of the following more general problem involving two matroids:
Double MC-Multiway-Cut
Input: A graph $G=(V,E)$, edge weights $w\in\mathbb{R}^{E}_{+}$, and rank-$k$ matroids $\mathcal{M}_{1}=(V,\mathcal{B}_{1})$ and $\mathcal{M}_{2}=(V,\mathcal{B}_{2})$ given by independence oracles, where $\mathcal{B}_{i}$ is the family of bases of $\mathcal{M}_{i}$.
[…]
Here, the goal is to find a minimum multiway cut where there is a choice of terminals forming a basis of $\mathcal{M}_{1}$, and not necessarily the same choice of terminals forming a basis of $\mathcal{M}_{2}$.
Theorem 11
Double MC-Multiway-Cut can be solved in polynomial time if $G$ is a tree.
Proof: We can construct $\mathcal{M}_{i}^{\prime}$ as in (1) using $\mathcal{M}_{i}$ ($i=1,2$), and we can find a minimum-weight common spanning set $C$ of $\mathcal{M}_{1}^{\prime}$ and $\mathcal{M}_{2}^{\prime}$ in polynomial time. Let $\{V_{1},\dots,V_{\ell}\}$ denote the connected components of $E\setminus C$, and let $B_{1}\in\mathcal{B}_{1}$ and $B_{2}\in\mathcal{B}_{2}$ be bases such that $|B_{j}\cap V_{i}|\leq 1$ for all $i\in[\ell],j\in[2]$. If $\ell>k$, then there must be classes $V_{i}$ and $V_{j}$ such that $|(V_{i}\cup V_{j})\cap B_{1}|\leq 1$ and $|(V_{i}\cup V_{j})\cap B_{2}|\leq 1$. Remove $V_{i}$ and $V_{j}$ from the partition, and add $V_{i}\cup V_{j}$. Repeat this operation until $\ell=k$, and let $\{V_{1}^{\prime},\dots,V_{k}^{\prime}\}$ be the obtained feasible partition. Then $\sum_{i=1}^{k}d_{w}(V_{i}^{\prime})\leq\sum_{i=1}^{k}d_{w}(V_{i})$, so $\{V_{1}^{\prime},\dots,V_{k}^{\prime}\}$ is an optimal solution, because the boundary of any feasible partition is a common spanning set of $\mathcal{M}_{1}^{\prime}$ and $\mathcal{M}_{2}^{\prime}$.  $\Box$
We get a significantly more difficult problem if the representatives for the two matroids are required to be the same, that is, if we require the set of representatives to form a common basis of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. The problem is defined as follows.
Common MC-Multiway-Cut
Input: A graph $G=(V,E)$, edge weights $w\in\mathbb{R}^{E}_{+}$, and rank-$k$ matroids $\mathcal{M}_{1}=(V,\mathcal{B}_{1})$ and $\mathcal{M}_{2}=(V,\mathcal{B}_{2})$ given by independence oracles, where $\mathcal{B}_{i}$ is the family of bases of $\mathcal{M}_{i}$.
Goal: Find a solution to $\min\left\{\sum_{i=1}^{k}d_{w}(V_{i})\colon\{V_{i}\}_{i=1}^{k}\text{ is a partition of }V\text{ s.t. there is }B\in\mathcal{B}_{1}\cap\mathcal{B}_{2}\text{ with }|B\cap V_{i}|=1\text{ for all }i\in[k]\right\}.$
We show that Common MC-Multiway-Cut includes the problem of finding a common basis in the intersection of three partition matroids; this latter problem is known to be hard as it generalizes e.g. the 3-dimensional matching problem.
Theorem 12
The problem of verifying whether three partition matroids have a common basis reduces to the problem of deciding if the optimum of a Common MC-Multiway-Cut instance is 00.
Proof: Let $P_{1},P_{2},P_{3}$ be partition matroids on a ground set $S$ with partition classes $\mathcal{U}=\{U_{1},\dots,U_{r}\}$, $\mathcal{V}=\{V_{1},\dots,V_{r}\}$, and $\mathcal{W}=\{W_{1},\dots,W_{r}\}$, respectively. We encode $P_{3}$ into a tree of depth two by creating a root $z$, a depth one vertex $v_{i}$ for each $W_{i}$, and leaf vertices for each element of $s\in S$. The edges are between $v_{i}$ and $z$, and for each $s\in W_{i}$, we add an edge between $s$ and $v_{i}$. We add the new vertices to the ground set of $P_{1}$ and $P_{2}$ to create $M_{1}$ and $M_{2}$: each $v_{i}$ is added as a loop in both matroids, while $z$ is added as a free element, i.e., it is added to every base. We assign weight 0 to the edges from the root $z$, and all other edges get unit weights.
The only possible solution with objective value 00 is the partition formed by $\{z\}$ and $W_{i}+v_{i}$ ($i\in[r]$). This is a feasible solution if and only if $P_{1}$ and $P_{2}$ have a common basis that intersects each $W_{i}$ in exactly one element, that is, $P_{1}$, $P_{2}$, and $P_{3}$ have a common basis.  $\Box$
3.4 Paving Matroids
In this section, we show that Sub-MCP is equivalent to Sub-k-P when the matroid $\mathcal{M}$ is a paving matroid. The key observation is that if the matroid is a rank-$k$ paving matroid, then any $k$-partition has a valid basis representative.
Lemma 13
Let $\mathcal{M}=(V,\mathcal{B})$ a rank-$k$ paving matroid for some $k\geq 1$ and $\mathcal{P}=\{V_{1},\dots,V_{k}\}$ be a partition of $V$. Then there exists $B\in\mathcal{B}$ such that $|B\cap V_{i}|=1$ for $i\in[k]$.
Proof: Let =$\mathcal{H}=\{H_{1},\dots,H_{q}\}$ be the hypergraph representation of $\mathcal{M}$. Since every set of size at most $k-1$ is independent in $\mathcal{M}$, there is an independent set $X=\{v_{1},\dots,v_{k-1}\}$ such that $v_{i}\in V_{i}$ for $i\in[k-1]$. If $X+v$ is a basis for some $v\in V_{k}$, then we are done. Otherwise, for every $v\in V_{k}$ there exists a hyperedge $H_{v}\in\mathcal{H}$ such that $X+v\subseteq H_{v}$. As $|H_{u}\cap H_{v}|=|X|=k-1$ for any distinct $u,v\in V_{k}$, we get that $H_{u}=H_{v}$ by the properties of the hypergraph representation of the paving matroid. Let us denote this unique hyperedge by $H$; we thus have $V_{k}\subseteq H$. By repeating the same argument for each partition class in place of $V_{k}$, we get that $H=V$, contradicting the hyperedges being proper subsets of $V$.  $\Box$
4 Greedy Splitting for Sub-MCP
In the analysis of the greedy splitting algorithm by Zhao, Nagamochi and Ibaraki [17], they rely on a key lemma in their approximation guarantees for Sub-MP, SymSub-MP, MonSub-MP, Sub-k-P, SymSub-k-P, and MonSub-k-P. We show that this lemma holds even when an additional matroid constraint is imposed, and then follow the same method to prove Theorems 2, 3, and 4; all aforementioned results of [17] follow from these three theorems as well. The modified greedy splitting algorithm is described as Algorithm 2.
Algorithm 2 Greedy Splitting Algorithm for Sub-MCP
1:Input: A submodular function $f\colon 2^{V}\to\mathbb{R}_{+}$ and a rank-$k$ matroid $\mathcal{M}=(V,\mathcal{I})$.
2:Output: A feasible partition $\mathcal{P}$.
3:$\mathcal{P}_{1}\leftarrow\{V\}$.
4:for $i=1,\dots,k-1$ do
[…]
6:     $\mathcal{P}_{i+1}\leftarrow(\mathcal{P}_{i}\setminus W)\cup\{X,W\setminus X\}$
7:end for
8:Return $\mathcal{P}_{k}$.
Remark 14
Line 3 of Algorithm 2 can be executed in polynomial time by the following subroutine.
  1. 1.
    For every pair $x,y$ of elements in $W$, we can decide if $x,y$ can be extended to an independent set so that we pick a single element from each class of $\mathcal{P}_{i}\setminus W$.
  2. 2.
    Now that we know which pairs $x,y$ are possible representatives, we can do the following: for each fixed pair $x,y$, find the best partition of $W$ into $X$ and $W\setminus X$ such that $x\in X$, $y\in W\setminus X$.
Lemma 15 (Main Lemma in [17])
Let $\mathcal{P}_{i}$ be the $i^{th}$ partition returned by the greedy splitting procedure, and $\mathcal{P}=(V_{1},\dots,V_{i})$ be any partition satisfying the matroid constraint. Then for any submodular $f$,
$f(\mathcal{P}_{i})\leq\sum_{j=1}^{i-1}(f(V_{j})+f(V\setminus V_{j}))-(i-2)f(V)$
Proof: We proceed, as in [17], by induction on $i$. The base case $i=1$ holds trivially. In the $i^{th}$ step, we are given a partition $\mathcal{P}=(V_{1},\dots,V_{i})$ along with some $I_{i}\in\mathcal{I}$ that is also independent in the partition matroid induced by $\mathcal{P}$. Let $\mathcal{P}_{i-1}=(U_{1},\dots,U_{i-1})$ be the partition created by the algorithm in the previous step, and $I_{i-1}\in\mathcal{I}$ be the representatives corresponding to $\mathcal{P}_{i-1}$. As $|I_{i}|\geq|I_{i-1}|$, there must be an element $v\in I_{i}\setminus I_{i-1}$ such that $I_{i-1}+v$ is independent in $\mathcal{M}$. The element $v$ must be in some part of $\mathcal{P}_{i-1}$, say $U_{h}$, and some part of $\mathcal{P}$, say $V_{j}$. Let $u\in U_{h}$ be the element of $I_{i-1}$ in $U_{h}$.
Case 1: Say $u\notin V_{j}\cap U_{h}$. Then $V_{j}\cap U_{h},U_{h}\setminus V_{j}$ is a valid candidate split for Line 4 in Algorithm 2, so
$f(\mathcal{P}_{i})-f(\mathcal{P}_{i-1})\leq f(V_{j}\cap U_{h})+f(U_{h}\setminus V_{j})-f(U_{h}).$
Submodularity then implies
$f(\mathcal{P}_{i})-f(\mathcal{P}_{i-1})\leq f(V_{j})+f(U_{h}\setminus V_{j})-f(V_{j}\cup U_{h})\leq f(V_{j})+f(V\setminus V_{j})-f(V).$
Case 2: For the other case, assume $u\in V_{j}\cap U_{h}$. Then $u\notin I_{i}$ since the element of $I_{i}$ in $V_{j}$ is $v$; furthermore, $I_{i-1}^{\prime}=I_{i-1}+v-u$ is a valid representative for $\mathcal{P}_{i-1}$. We can repeat the same argument with $I_{i-1}^{\prime}$. This process will eventually terminate, as $|I_{i-1}^{\prime}\cap I_{i}|=|I_{i-1}\cap I_{i}|+1$, hence after a finite number of steps either we are in Case 1, or $I_{i-1}\subseteq I_{i}$. In the latter case, $u$ cannot be in $V_{j}$, as there is another $v\in V_{j}\cap I_{i}$, and we are again in Case 1.
Thus, there is some $V_{j}\in\mathcal{P},$ such that $f(\mathcal{P}_{i})-f(\mathcal{P}_{i-1})\leq f(V_{j})+f(V\setminus V_{j})-f(V)$. By the induction hypothesis, when applied to $\mathcal{P}\setminus V_{j}$, we get
$f(\mathcal{P}_{i-1})\leq\sum_{j=1}^{i-2}(f(V_{j})+f(V\setminus V_{j}))-(i-3)f(V).$
Therefore,
$f(\mathcal{P}_{i})\leq f(\mathcal{P}_{i-1})+f(V_{j})+f(V\setminus V_{j})-f(V)\leq\sum_{j=1}^{i-1}(f(V_{j})+f(V\setminus V_{j}))-(i-2)f(V),$
concluding the proof of the lemma.  $\Box$
With Lemma 15 in hand, the results of [17] follow immediately. In order to make the paper self-contained, we repeat them here.
Proof: [Proof of Theorem 4] Lemma 15 shows that for the optimal partition $\mathcal{P}=\{V_{1},\dots,V_{k}\}$, the output $\mathcal{P}_{k}$ of the Algorithm 2 satisfies
$f(\mathcal{P}_{k})\leq\sum_{j=1}^{k-1}(f(V_{j})+f(V\setminus V_{j}))-(k-2)f(V)\leq\sum_{j=1}^{k-1}\sum_{i=1}^{k}f(V_{i})-(k-2)f(V)\leq(k-1)f(\mathcal{P}).$
This concludes the proof.  $\Box$
Proof: [Proof of Theorem 2] Let $\mathcal{P}^{*}=\{V_{1}^{*},\dots,V_{k}^{*}\}$ be the optimal partition for SymSub-MCP, ordered so that $V_{1}^{*}\leq V_{2}^{*}\leq\dots\leq V_{k}^{*}$. Then by Lemma 15,
$f(\mathcal{P}_{k})\leq\sum_{j=1}^{k-1}(f(V_{j}^{*})+f(V\setminus V_{j}^{*}))-(k-2)f(V),$
where $\mathcal{P}_{k}$ is the partition returned by Algorithm 2. Using the fact that $f$ is symmetric and submodular, we have
$f(\mathcal{P}_{k})\leq 2\sum_{j=1}^{k-1}f(V_{j}^{*})-(k-2)f(V)\leq 2\sum_{j=1}^{k-1}f(V_{j}^{*})\leq\left(2-\frac{2}{k}\right)f(\mathcal{P}^{*}),$
concluding the proof of the theorem.  $\Box$
Proof: [Proof of Theorem 3] Let $\mathcal{P}^{*}=\{V_{1}^{*},\dots,V_{k}^{*}\}$ be the optimal partition for MonSub-MCP, ordered so that $V_{1}^{*}\leq V_{2}^{*}\leq\dots\leq V_{k}^{*}$. Then, by Lemma 15,
$f(\mathcal{P}_{k})\leq\sum_{j=1}^{k-1}(f(V_{j}^{*})+f(V\setminus V_{j}^{*}))-(k-2)f(V),$
where $\mathcal{P}_{k}$ is the partition returned by Algorithm 2. Using the fact that $f$ is monotone and submodular, we have
$f(\mathcal{P}_{k})$ $\leq\sum_{j=1}^{k-1}f(V_{j}^{*})+f(V\setminus V_{k-1}^{*})\leq\sum_{j=1}^{k-2}f(V_{j}^{*})+f(V_{k-1}^{*})+f(V\setminus V_{k-1}^{*})$
$\leq\left(1-\frac{2}{k}\right)f(\mathcal{P}^{*})+f(V_{k-1}^{*})+f(V\setminus V_{k-1}^{*})\leq\left(2-\frac{2}{k}\right)f(\mathcal{P}^{*}),$
concluding the proof of the theorem.  $\Box$
Our analysis of the $(2-2/k)$-approximation for monotone submodular functions is tight for the greedy splitting algorithm, even without the matroid constraints, by the following lemma.
Lemma 16
There is an instance of monotone submodular $k$-partition for which the greedy splitting algorithm achieves exactly $(2-2/k)$-approximation.
Proof: Let $\mathcal{P}=\{S_{1},\dots,S_{k}\}$ be a partition of the ground set $V$ into subsets of size at least $2$. Consider the laminar matroid $\mathcal{M}$ in which a set $X\subseteq V$ is independent if and only if $|V\cap S_{i}|\leq 1$ for $i\in[k]$ and $|X|\leq k-1$. Let $f$ be the rank function of $\mathcal{M}$ – clearly, $f$ is monotone and submodular.
For any initial split of the greedy splitting algorithm, we have $f(S)+f(V\setminus S)\geq k$. Therefore, the algorithm might choose $S$ to be a singleton. In general, as the greedy splitting algorithm proceeds, in the $i^{th}$ step the algorithm might split the class $S_{i}$ into a singleton and the set of remaining elements. In such a scenario, the resulting partition has the form $\{s_{1},\dots,s_{k-1},V\setminus\{s_{1},\dots,s_{k-1}\}\}$. The sum of the ranks of these sets is $22k-$ 2, while the optimal partition is $\mathcal{P}$ with a total $f$ value of $k$. This shows that greedy splitting might lead to a multiplicative error of $2-2/k$.  $\Box$